# Enhancing Quantum Memory Lifetime with Measurement-Free Local Error Correction and Reinforcement Learning

### v1

@@ I Introduction
In this work, we optimize circuit-level LEC protocols that consist of faulty multi-qubit gates — to perform both syndrome extraction and ancilla-controlled error removal — by developing an RL framework. In particular, our RL framework takes a fixed set of faulty gates as inputs and outputs an LEC circuit, with optimized length and layout. By performing extensive circuit-level simulations, we show these circuits successfully reduce the error density when gate fidelities are high enough. We then characterize the resulting extension in the lifetime of a memory encoded in three finite-sized systems: 2D toric code, 2D Ising model, and 4D toric code. Finally, we explore two potential applications of this scheme: (1) reducing the rate of mid-circuit readouts required to preserve quantum memory and (2) dissipatively preparing quantum states of topological phases.
This paper is organized as follows: we begin in Section II by outlining the key insights and main results of this work. We provide a detailed analysis of a circuit-level LEC scheme in Section III. More specifically, we present an implementation of the basic components of the LEC circuits (see Section III.1) followed by characterization and resolution of the error patterns that limit conventional LEC circuits (see Section III.2 and III.3). Then, we discuss a setting for the RL framework in Section IV.1 and its optimization of circuit depth and layout within each code in Section IV.2. We discuss a quantitative characterization of memory lifetime extension by LEC circuits and a comparison between RL-optimized and conventional LEC circuits in Section V. Furthermore, following the considerations of the experimental implementation of the LEC scheme in Section VI.1, we propose two potential applications of this scheme: reducing the rate of mid-circuit readouts required to preserve quantum memory in 2D toric code (see Section VI.2) and dissipatively preparing quantum states of topological phases (see Section VI.3). Finally, we present conclusions and outlook in Section VII.
@@ II Overview of Main Results
We focus on a circuit-level, measurement-free LEC protocol, composed of faulty multi-qubit gates. As illustrated in Fig. 1, the circuits are constructed from two types of basic local quantum operations: (1) ancilla-based syndrome extraction with a series of two-qubit gates and (2) ancilla-controlled error removal with three-qubit gates. We focus on three finite-sized error-correcting codes — the 2D toric code, classical 2D Ising model, and 4D toric code — and start by studying the performance of existing LEC approaches, such as nearest-neighbor matching in 2D toric code [46] or Toom’s rule in 2D Ising model and 4D toric code [38]. We next develop and apply a new RL framework to optimize LEC circuits and improve the overall performance in all three cases. Finally, we provide a detailed and intuitive understanding of how the optimization achieves improved performance.
@@
Figure 1: Overview of circuit-level local error correction (LEC) protocol and the reinforcement learning (RL) optimization framework. The LEC circuit is a sequence of local coherent quantum operations — not including a mid-circuit projective readout of any qubit — applied to data qubits (circle) and ancillas (triangle). It takes data qubits with errors (colored red) and initialized ancillas as input and moves errors of data qubits into ancillas. This local decoding can be repeated over multiple cycles with a finite number of ancillas by resetting them into the initialized ones. A reinforcement learning (RL) framework constructs an optimized LEC circuit by combining two types of “building block” operations, followed by corresponding coherent and independent reconfiguration of qubits. First, ancilla-based syndrome extraction operations enable the ancillas to measure and store syndrome information that indicates the existence of errors on neighboring data qubits. On the other hand, ancilla-controlled error removal operations provide local feedback operation controlled by ancillas onto data qubits to correct their errors. All these operations can be achieved experimentally within the state-of-the-art quantum processors, including the Rydberg atom arrays.
@@
We find that the existing LEC methods are limited by specific low-weight error patterns, which are not removed by the local error removal operations comprising the circuit, as also noted by previous literature [44, 40] (see Section III.2). To address this, we introduce additional longer-range operations that are capable of removing some of these error patterns and thus increase the weight of the minimum uncorrectable error in each code (see Section III.3). The RL framework, as shown in Fig. 6, assembles the operations into optimized LEC circuits that prevent the accumulation of higher-weight error patterns on average over multiple rounds of their application. In general, since we are considering a circuit-level noise model, adding additional gates also increases the number of errors. Importantly, we find that the RL can balance this cost, with the benefits arising from removing such errors: as gates get noisier, the optimization prefers shorter and simpler LEC circuits to minimize errors. In contrast, when we consider gates with higher fidelity, the RL procedure uses more complex decoding procedures to further improve performance (see Section IV.2).
@@
We illustrate the utility of RL-optimized LEC circuits by showing how they can extend the lifetime of quantum memory in a sub-threshold gate error regime. In each LEC round, we introduce “ambient” errors, which model noises arising from other faulty operations including idling or transversal logical gates [6, 71], with $p_{amb}$ error rate. Then, we apply a full LEC circuit that consists of faulty multi-qubit controlled gates with $p_{gate}$ error rate. We define an average logical qubit lifetime $T$ to be an average number of LEC rounds before encountering the first failure in decoding. Simulating different values of $p_{amb}$, we observe a scaling of the error suppression consistent with a power law dependence upon $T$ and $p_{amb}$ (see Fig. 10). To quantify this effect, we introduce the notion of an effective code distance $D_{eff}$ as a function of a given LEC circuit and $p_{gate}$ such that $T\propto(p_{amb})^{-D_{eff}}$. $D_{eff}$ is related to the ability of a noisy decoding circuit to remove (on average) error patterns of a certain weight (see Section V.2). In the sub-threshold $p_{gate}$ regime, we numerically show that RL-optimized LEC circuits have higher $D_{eff}$ compared to conventional LEC circuits in each code and are thus more effective in removing errors from noisy states (see Fig. 11).
@@ III Circuit-level LEC Scheme
In this section, we overview our family of circuit-level LEC schemes in 2D toric code, classical 2D Ising model, and 4D toric code. Here, we focus on providing an explicit translation into a quantum circuit with multi-qubit gates from the local decoding rule presented in previous literature [74, 75, 38, 46]. Section III.1 discusses how to implement the basic components of the LEC circuit. Section III.2 characterizes the error patterns that limit conventional LEC circuits. Then, Section III.3 introduces the hierarchically extended set of error removal operations, to address this limitation of the circuit-level LEC scheme.
@@ III.1 Circuit-level LEC operations in each code
We will use three models to develop and test our framework, illustrated in Fig. 2. The 2D toric code is a canonical quantum error-correcting code, which has been studied extensively [9, 10]. This code, having point-like syndrome excitations, cannot scalably protect quantum information with a local decoding protocol. Nevertheless, we still observe pseudo-threshold behavior, as discussed further below, and gain simple intuition for how LEC works using this model. To extend our analysis to models that can, in principle, be protected by an LEC procedure, we consider the 2D (classical) Ising model, which realizes a classical memory [76], and the 4D toric code, which realizes a quantum memory [9, 77]. Both have line-like syndrome excitations, which can be corrected via local operations [38, 74, 75, 78]. With sufficiently low error rates, LEC circuits can protect their information indefinitely, making these codes self-correcting memories. Each model can be defined by its stabilizers, where their eigenvalues are either $+1$ or $-1$. Then, each model encodes logical qubit(s) where all stabilizers have $+1$ eigenvalues.
  • 2D toric code [79]:
    As in Fig. 2(a), 2D toric code can be represented on the 2D lattice with periodic boundary conditions. Data qubits live at edges of this lattice, and there are plaquette-type ($A_{p}$) and vertex-type ($B_{v}$) stabilizers defined as a tensor product of four Pauli operators on data qubits:
    $\hat{A}_{p}=\bigotimes_{e\in p}\hat{Z}_{e},\quad\hat{B}_{v}=\bigotimes_{e\in v}\hat{X}_{e}$ (1)
    where $e$ corresponds to the index of an edge neighboring the vertex $v$ or the plaquette $p$. Note that a 2D toric code lattice encodes two logical qubits with logical $\hat{Z}$ (or $\hat{X}$) that correspond to tensor products of $\hat{Z}$ (or $\hat{X}$) operators along the topologically non-trivial loops of the 2D torus.
  • 2D Ising model [9]:
    As in Fig. 2(b), the 2D Ising model can also be represented on the 2D lattice with periodic boundary conditions. Data qubits live on plaquettes of this lattice, and there are vertical-type ($A_{v}$) and horizontal-type ($A_{h}$) stabilizers that are tensor products of two $\hat{Z}$ operators on the data qubits:
    $\hat{A}_{v,(i,j)}$ $=\hat{Z}_{(i,j)}\otimes\hat{Z}_{(i,j+1)},$ (2)
    $\hat{A}_{h,(i,j)}$ $=\hat{Z}_{(i,j)}\otimes\hat{Z}_{(i+1,j)}$ (3)
    where $(i,j)$ denotes a 2D coordinate of a plaquette. Since the 2D Ising model can detect only a bit-flip error on the qubit, it encodes a classical memory. Also, the entire 2D Ising model lattice encodes one logical qubit that consists of all qubits with either |$\ket{0}$ or |$\ket{1}$.
  • 4D toric code [9]:
    As in Fig. 2(c), 4D toric code can be represented on the 4D lattice with periodic boundary conditions. Data qubits live on the faces of this lattice, and there are edge-type ($A_{e}$) and cube-type ($B_{c}$) stabilizers. Each stabilizer is a tensor product of six Pauli operators on data qubits:
    $\hat{A}_{e}=\bigotimes_{f\in e}\hat{Z}_{e},\quad\hat{B}_{c}=\bigotimes_{f\in c}\hat{X}_{c},$ (4)
    where $f$ corresponds to the index of a face neighboring the edge $e$ or the cube $c$. Note that a 4D toric code lattice encodes two logical qubits with logical $\hat{Z}$ (or $\hat{X}$) that correspond to tensor products of $\hat{Z}$ (or $\hat{X}$) operators on the topologically non-trivial sheets of the periodic 4D lattice.
@@
Figure 3: The gate-level implementation of syndrome extraction and error removal operation within each error-correcting code. (a) Syndrome extraction operation using ancillas is a sequence of CNOT gates in parallel. An example of such a sequence with unit cells for 2D toric code is shown. To store eigenvalues of stabilizers that consist of $\hat{Z}$s in ancillas, applied CNOT gates are controlled by data qubits and targeted on ancillas, and vice versa for stabilizers that consist of $\hat{X}$s, following the optimized order as discussed in Appendix A. (b) Error removal operation to reduce error density of data qubits controlled by ancillas is a single application of CCX or CCZ gates in parallel. An example of the simplest such operation with a unit cell for each code is shown. To remove $\hat{X}$ (or $\hat{Z}$) errors, CCX (or CCZ) gates are applied.
@@
The syndrome extraction circuits are used to measure stabilizers via ancillas. For each of the above error-correcting codes, the stabilizers are a tensor product of $\hat{Z}$ or $\hat{X}$ operators on the data qubits. Since $\hat{Z}$ and $\hat{X}$ operators anti-commute with each other, eigenvalue $-1$ of stabilizer $\hat{S}_{Z}$ (or $\hat{S}_{X}$) denotes that there exist odd number of $\hat{X}$ (or $\hat{Z}$) errors on data qubits that support the stabilizer. Similarly, eigenvalue $+1$ of $\hat{S}_{Z}$ (or $\hat{S}_{X}$) denotes that there exists an even number of $\hat{X}$ (or $\hat{Z}$) errors on data qubits that support the stabilizer. This provides a simple procedure to extract syndromes. For 2D toric code, we consider an ancilla initialized to |$\ket{0}$ for every plaquette. Then, we apply a sequence of four CNOT gates controlled by data qubits at edges neighboring the plaquette and targeted on the ancilla. Due to the commutation relation between a CNOT gate and $\hat{X}$ and $\hat{Z}$ operator, we obtain
$\hat{X}^{(1+S_{Z})/2}\ket{0}=\begin{cases}\ket{0}&\text{if }S_{Z}=+1\\ \ket{1}&\text{if }S_{Z}=-1\end{cases},$ (5)
where $S_{Z}$ is an eigenvalue of $\hat{S}_{Z}$. Similarly, using two Hadamard gates and a sequence of four CNOT gates controlled by the ancilla and target on data qubits, we obtain ancillas on vertices storing eigenvalues of $\hat{S}_{X}$. For a more efficient syndrome extraction operation, we parallelize the CNOT gates to the entire 2D toric code lattice as shown in Fig. 3(a). Similar sequences of CNOT gates in parallel can detect errors by syndrome extraction for the 2D Ising model and 4D toric code. Note that we assume a reset of ancillas to their initial states |$\ket{0}$ before a new syndrome extraction operation.
The error removal operations are more complex and involve applying correction gates conditioned on the state of the ancillas. Let |$\ket{A_{1}}$ and |$\ket{A_{2}}$ be two ancillas that store an eigenvalue of stabilizer $\hat{S}_{Z,A_{1}}$ and $\hat{S}_{Z,A_{2}}$, respectively. Then, applying CCX gate controlled by these two ancillas — without knowing their quantum states — and targeted on the data qubit |$\ket{D}$ results in
[…]
@@
As the simplest scenario, consider $\hat{S}_{Z,A_{1}}$ and $\hat{S}_{Z,A_{2}}$ share one supporting data qubit |$\ket{D}$. If $S_{Z,A_{1}}=S_{Z,A_{2}}=-1$, i.e. |$\ket{A_{1}}=\ket{A_{2}}=\ket{1}$, there are odd number of errors on data qubits supporting $\hat{S}_{Z,A_{1}}$ and in $\hat{S}_{Z,A_{2}}$ each. When data qubits have low error density, it is likely that the shared data qubit |$\ket{D}$ has an $\hat{X}$ error, which gets corrected by applying the above CCX gate. We can similarly remove a $\hat{Z}$ error on the data qubit |$\ket{D}$ by applying CCZ gate controlled by ancillas |$\ket{A_{3}}$ and |$\ket{A_{4}}$, storing an eigenvalue of stabilizer $\hat{S}_{X,A_{3}}$ and $\hat{S}_{X,A_{4}}$ respectively, and targeted on |$\ket{D}$. The simplest gate layout for an error removal operation in each code is shown in Fig. 3(b).
@@
A typical global decoding approach also uses ancilla-based syndrome extraction operations. However, instead of performing the error removal operation via local three-qubit gates, all the ancillas are projectively measured after the syndrome extraction operation. Then, a classical decoding algorithm globally processes readout outcomes to infer the most likely underlying errors, which can be subsequently corrected by applying local Pauli operators. However, in our LEC scheme, only certain local correction operations are implemented, leading to the existence of additional uncorrectable error configurations (see Section III.2).
We explain how we implement a code base for circuit-level error correction benchmark simulation for each error-correcting code in more detail in Appendix A.
III.2 Simplest LEC operation and its limitation

### v2

@@ I Introduction
In this work, we optimize circuit-level LEC protocols that consist of faulty multi-qubit gates — to perform both syndrome extraction and ancilla-controlled error removal — by developing an RL framework. In particular, our RL framework takes a fixed set of faulty gates as inputs and outputs an LEC circuit, with optimized length and layout. By performing extensive circuit-level simulations, we show these circuits successfully reduce the error density when gate fidelities are high enough. We then characterize the resulting extension in the lifetime of a memory encoded in three finite-sized systems: 2D toric code, 2D Ising model, and 4D toric code. Finally, we explore two near-term applications of this scheme: (1) reducing the rate of mid-circuit readouts required to preserve quantum memory and (2) dissipatively preparing quantum states of topological phases.
This paper is organized as follows: we begin in Section II by outlining the key insights and main results of this work. Section III provides a detailed analysis of a circuit-level LEC scheme. In particular, we present an implementation of the basic components of the LEC circuits followed by characterization and resolution of the error patterns that limit conventional LEC circuits. Then, Section IV discusses a setting for the RL framework and its optimization of circuit depth and layout within each code. Section V presents a quantitative characterization of memory lifetime extension by LEC circuits and a comparison between RL-optimized and conventional LEC circuits. Furthermore, Section VI proposes two near-term applications of this scheme: reducing the rate of mid-circuit readouts required to preserve quantum memory in 2D toric code and dissipatively preparing quantum states of topological phases. Finally, we present conclusions and outlook in Section VII.
@@ II Overview of Main Results
We focus on a circuit-level, measurement-free LEC protocol, composed of faulty multi-qubit gates. As illustrated in Fig. 1, the circuits are constructed from two types of basic local quantum operations: (1) ancilla-based syndrome extraction with a series of two-qubit gates and (2) ancilla-controlled error removal with three-qubit gates. We focus on three finite-sized error-correcting codes — the 2D toric code, classical 2D Ising model, and 4D toric code — and start by studying the performance of existing LEC approaches, such as nearest-neighbor matching in 2D toric code [46] or Toom’s rule in 2D Ising model and 4D toric code [38]. We then develop and apply a new RL framework to optimize LEC circuits and improve the overall performance in all three cases. Finally, we provide a detailed and intuitive understanding of how the optimization achieves improved performance.
@@
Figure 1: Overview of circuit-level local error correction (LEC) protocol and the reinforcement learning (RL) optimization framework. The LEC circuit is a sequence of local coherent quantum operations — not including a mid-circuit projective readout of any qubit — applied to data qubits (circle) and ancillas (triangle). It takes data qubits with errors (colored red) and initialized ancillas as input and moves errors of data qubits into ancillas. This local decoding can be repeated over multiple cycles with a finite number of ancillas by resetting them into the initialized ones. A reinforcement learning (RL) framework constructs an optimized LEC circuit by combining two types of “building block” operations, followed by corresponding coherent and independent reconfiguration of qubits. First, ancilla-based syndrome extraction operations enable the ancillas to measure and store syndrome information that indicates the existence of errors on neighboring data qubits. On the other hand, ancilla-controlled error removal operations provide local feedback operation controlled by ancillas onto data qubits to correct their errors. All these operations can be achieved experimentally within the state-of-the-art quantum processors, including the Rydberg atom arrays (see Appendix A).
@@
We find that the existing LEC methods are limited by specific low-weight error patterns, which are not removed by the local error removal operations comprising the circuit, as also noted by previous literature [44, 40]. To address this, we introduce additional longer-range operations that are capable of removing some of these error patterns and thus increase the weight of the minimum uncorrectable error in each code. The RL framework, as shown in Fig. 6, assembles the operations into optimized LEC circuits that prevent the accumulation of higher-weight error patterns on average over multiple rounds of their application. In general, since we are considering a circuit-level noise model, adding additional gates also increases the number of errors. Importantly, we find that the RL can balance this cost, with the benefits arising from removing such errors: as gates get noisier, the optimization prefers shorter and simpler LEC circuits to minimize errors. In contrast, when we consider gates with higher fidelity, the RL procedure uses more complex decoding procedures to further improve performance.
@@
We illustrate the utility of RL-optimized LEC circuits by showing how they can extend the lifetime of quantum memory in a sub-threshold gate error regime. In each LEC round, we introduce “ambient” errors, which model noises arising from other faulty operations including idling or transversal logical gates [6, 71], with $p_{\rm{amb}}$ error rate. Then, we apply a full LEC circuit that consists of faulty multi-qubit controlled gates with $p_{\rm{gate}}$ error rate. We define an average logical qubit lifetime $T$ to be an average number of LEC rounds before encountering the first failure in decoding. Simulating different values of $p_{\rm{amb}}$, we observe a scaling of the error suppression consistent with a power law dependence upon $T$ and $p_{\rm{amb}}$ (see Fig. 9). To quantify this effect, we introduce the notion of an effective code distance $D_{\rm{eff}}$ as a function of a given LEC circuit and $p_{\rm{gate}}$ such that $T\propto(p_{\rm{amb}})^{-D_{\rm{eff}}}$. $D_{\rm{eff}}$ is related to the ability of a noisy decoding circuit to remove (on average) error patterns of a certain weight. In the sub-threshold $p_{\rm{gate}}$ regime, we numerically show that RL-optimized LEC circuits have higher $D_{\rm{eff}}$ compared to conventional LEC circuits in each code and are thus more effective in removing errors from noisy states (see Fig. 10).
@@ III Circuit-level LEC Scheme
In this section, we overview our family of circuit-level LEC schemes in 2D toric code, classical 2D Ising model, and 4D toric code by providing an explicit translation into a quantum circuit with multi-qubit gates from the local decoding rule [74, 75, 38, 46]. Section III.1 discusses how to implement the basic components of the LEC circuit. Section III.2 characterizes the error patterns that limit conventional LEC circuits and then introduces the hierarchically extended set of error removal operations to address this limitation of the circuit-level LEC scheme. To provide relevant details, we include the considerations of the experimental implementation of the LEC scheme in Appendix A and implementation of a code base outlined in this section in Appendix B.
@@ III.1 Circuit-level LEC operations in each code
We will use three models to develop and test our framework, illustrated in Fig. 2. The 2D toric code is a canonical quantum error-correcting code, which has been studied extensively [9, 10]. This code, having point-like syndrome excitations, cannot scalably protect quantum information with a local decoding protocol. Nevertheless, we still observe pseudo-threshold behavior, as discussed further below, and gain simple intuition for how LEC works using this model. To extend our analysis to models that can, in principle, be protected by an LEC procedure, we consider the 2D (classical) Ising model, which realizes a classical memory [76], and the 4D toric code, which realizes a quantum memory [9, 77]. Both have line-like syndrome excitations, which can be corrected via local operations [38, 74, 75, 78]. With sufficiently low error rates, LEC circuits can protect their information indefinitely, making these codes self-correcting memories. Each model can be defined by its stabilizers, where their eigenvalues are either $+1$ or $-1$. Then, each model encodes logical qubit(s) where all stabilizers have $+1$ eigenvalues. Note that each lattice of the error-correcting code below has a periodic boundary condition.
  • 2D toric code [79]:
    As in Fig. 2(a), data qubits live on edges of the 2D lattice, and plaquette-type ($\hat{A}_{p}$) and vertex-type ($\hat{B}_{v}$) stabilizers are tensor products of four Pauli $\hat{Z}$s (or $\hat{X}$s) on data qubits neighboring each plaquette $p$ (or vertex $v$). Two logical qubits are encoded with logical $\hat{Z}$ (or $\hat{X}$) that correspond to tensor products of $\hat{Z}$s (or $\hat{X}$s) along the topologically non-trivial loops.
  • 2D Ising model [9]:
    As in Fig. 2(b), data qubits live on plaquettes of the 2D lattice, and vertical-type ($\hat{A}_{v}$) and horizontal-type ($\hat{A}_{h}$) stabilizers are tensor products of two $\hat{Z}$s on the data qubits adjacent to each vertical or horizontal edge. Since the 2D Ising model can detect only a bit-flip error on the qubit, it encodes a classical memory. Also, one logical qubit is encoded that consists of all qubits with either |$\ket{0}$ or |$\ket{1}$.
  • 4D toric code [9]:
    As in Fig. 2(c), data qubits live on the faces of the 4D lattice, and there are edge-type ($\hat{A}_{e}$) and cube-type ($\hat{B}_{c}$) stabilizers. Each stabilizer is a tensor product of six $\hat{Z}$s (or $\hat{X}$s) on data qubits neighboring each edge $e$ (or cube $c$). Two logical qubits are encoded with logical $\hat{Z}$ (or $\hat{X}$) that correspond to tensor products of $\hat{Z}$s (or $\hat{X}$s) on the topologically non-trivial sheets.
@@
Figure 3: The gate-level implementation of syndrome extraction and error removal operation in each code. (a) Syndrome extraction operation using ancillas is a sequence of CNOT gates in parallel. An example of such a sequence with unit cells for 2D toric code is shown. Note that the CNOT gates are applied in the optimized order (see Appendix B). (b) Error removal operation to reduce error density of data qubits controlled by ancillas is a single application of CCX or CCZ gates in parallel. An example of the simplest such operation with a unit cell for each code is shown. To remove $\hat{X}$ (or $\hat{Z}$) errors, CCX (or CCZ) gates are applied.
@@
The syndrome extraction circuits are used to measure stabilizers via ancillas. For each of the above error-correcting codes, the stabilizers are a tensor product of $\hat{Z}$s or $\hat{X}$s on the data qubits. Since $\hat{Z}$ and $\hat{X}$ anti-commute with each other, eigenvalue of stabilizer $\hat{S}_{Z}$ (or $\hat{S}_{X}$) is $(-1)^{n_{s}}$ where $n_{s}$ is a number of $\hat{X}$ (or $\hat{Z}$) errors on data qubits that support the stabilizer. Due to the commutation relation between a CNOT gate and $\hat{X}$ and $\hat{Z}$, a sequence of CNOT gates performs a syndrome extraction. For example, for 2D toric code, we put an ancilla initialized to |$\ket{0}$ on every plaquette and then apply the sequence of parallelized CNOT gates as shown in Fig. 3(a). Then, each ancilla remains as |$\ket{0}$ if an eigenvalue of corresponding $\hat{S}_{Z}$ is +1 and flips into |$\ket{1}$ otherwise. Similar sequences of CNOT gates in parallel can detect errors by syndrome extraction for the 2D Ising model and 4D toric code. Note that we assume a reset of ancillas to their initial states |$\ket{0}$ before a new syndrome extraction operation.
@@
The error removal operations are more complex and involve applying correction gates conditioned on the state of the ancillas. Let |$\ket{A_{1}}$ and |$\ket{A_{2}}$ be two ancillas that store an eigenvalue of stabilizer $\hat{S}_{Z,A_{1}}$ and $\hat{S}_{Z,A_{2}}$, respectively. Then, applying CCX gate controlled by these two ancillas — without knowing their quantum states — and targeted on the data qubit |$\ket{D}$ results in $\hat{X}\ket{D}$ if both of the controlled qubits are |$\ket{1}$s and |$\ket{D}$ otherwise. As the simplest scenario, consider these two stabilizers share one supporting data qubit |$\ket{D}$. If both stabilizers have $-1$ eigenvalues, we have |$\ket{A_{1}}=\ket{A_{2}}=\ket{1}$ by syndrome extraction and there are an odd number of errors on data qubits supporting each of these stabilizers. When data qubits have low error density, it is likely that the shared data qubit |$\ket{D}$ has an $\hat{X}$ error, which gets corrected by applying the above CCX gate. We can similarly remove a $\hat{Z}$ error on the data qubit by applying a CCZ gate. The simplest gate layout for an error removal operation in each code is shown in Fig. 3(b).
@@
A typical global decoding approach also uses ancilla-based syndrome extraction operations. However, instead of performing the error removal operation via local three-qubit gates, all the ancillas are projectively measured after the syndrome extraction operation. Then, by globally processing readout outcomes, a classical decoding algorithm infers the most likely underlying errors, which can be subsequently corrected by applying local Pauli operators. However, in our LEC scheme, only certain local correction operations are implemented, leading to the existence of additional uncorrectable error configurations.
III.2 Extended LEC operations